# The theory of implicit operations

### v1

@@ 13. Absolutely closed and primal algebras
Let $\bm{A}$ be an L{L}-algebra, $\mathcal{F}\subseteq\mathsf{imp}_{pp}(\bm{A})$, and ${L}_{\mathcal{F}}$ an $\mathcal{F}$-expansion of L{L} such that $\bm{A}[{L}_{\mathcal{F}}]$ is defined. If $\bm{A}[{L}_{\mathcal{F}}]$ is primal, then $\mathbb{V}(\bm{A}[{L}_{\mathcal{F}}])$ is a Beth companion of $\mathbb{Q}(\bm{A})$. If, moreover, $\mathcal{F}\subseteq\mathsf{imp}_{eq}(\bm{A})$, then the Beth companion $\mathbb{V}(\bm{A}[{L}_{\mathcal{F}}])$ is equational.
@@
Suppose that $\bm{A}[{L}_{\mathcal{F}}]$ is primal. Then $\mathbb{V}(\bm{A}[{L}_{\mathcal{F}}])$ has the strong epimorphism surjectivity property by 13.24. Furthermore, recall from [21, Thm. IV.9.4] that $\mathbb{Q}(\bm{A}[{L}_{\mathcal{F}}])=\mathbb{V}(\bm{A}[{L}_{\mathcal{F}}])$ because $\bm{A}[{L}_{\mathcal{F}}]$ is primal. Together with 11.6, this yields that, to conclude that $\mathbb{V}(\bm{A}[{L}_{\mathcal{F}}])$ is a Beth companion of $\mathbb{Q}(\bm{A})$, it only remains to show that $\mathbb{Q}(\bm{A}[{L}_{\mathcal{F}}])$ is a pp expansion of $\mathbb{Q}(\bm{A})$.
Since $\mathcal{F}\subseteq\mathsf{imp}_{pp}(\bm{A})$, for every $f\in\mathcal{F}$ there exists a pp formula $\varphi_{f}$ functional in $\bm{A}$ that defines $f$. By 3.11 each $\varphi_{f}$ defines some $f^{*}\in\mathsf{imp}_{pp}(\mathbb{Q}(\bm{A}))$. Let $\mathcal{F}^{*}=\{f^{*}:f\in\mathcal{F}\}$. As $\bm{A}[{L}_{\mathcal{F}}]$ is defined, $f^{\bm{A}}$ is total for every $f\in\mathcal{F}$. Consequently, $f^{\bm{A}}=(f^{*})^{\bm{A}}$ because $f$ and $f^{*}$ are both defined by $\varphi_{f}$. Therefore, $(f^{*})^{\bm{A}}$ is total for every $f^{*}\in\mathcal{F}^{*}$. Then 8.11(ii) yields $\mathcal{F}^{*}\subseteq\mathsf{ext}_{pp}(\mathbb{Q}(\bm{A}))$. We can regard ${L}_{\mathcal{F}}$ as an $\mathcal{F}^{*}$-expansion of L{L} by setting $g_{f^{*}}=g_{f}$ for each $f\in\mathcal{F}$. Since $f^{\bm{A}}=(f^{*})^{\bm{A}}$ for every $f\in\mathcal{F}$, the definition of $\bm{A}[{L}_{\mathcal{F}}]$ is independent on whether ${L}_{\mathcal{F}}$ is thought of as an $\mathcal{F}$-expansion or as an $\mathcal{F}^{*}$-expansion. Thus, 10.5 implies that $\mathbb{Q}(\bm{A}[{L}_{\mathcal{F}}])$ is a pp expansion of $\mathbb{Q}(\bm{A})$ induced by $\mathcal{F}^{*}$ and ${L}_{\mathcal{F}}$.
@@
For $n\in\mathbb{Z}^{+}$, we denote by $\L _{n}$ the subalgebra of the real unit interval [0,1]{[0,1]} (cf. 13.17) with universe $\{\frac{m}{n}:m\in\mathbb{N},\ m\leqslant n\}$. Notice that $\L _{n}$ is a finite $\mathsf{MV}$-algebra of $n+1$ elements. We consider the variety
@@
As $\L _{n}\leqslant{[0,1]}$, 13.19(i, ii) implies that for all $a,b\in\L _{n}$,
@@
Therefore, $\psi_{n}$ defines a total unary $c_{n}\in\mathsf{imp}_{eq}(\L _{n})$ such that $c_{n}^{\scriptsize\L _{n}}$ is the constant map with value $\frac{1}{n}$. Let ${L}_{n}$ be a $c_{n}$-expansion of ${L}_{\mathsf{MV}_{n}}$. Since $c_{n}^{\scriptsize\L _{n}}$ is total, the algebra $\L _{n}[{L}_{n}]$ is defined and can thought of as the result of adding a constant for the element $\frac{1}{n}$ to $\L _{n}$. Then let $\mathsf{DMV}_{n}=\mathbb{V}(\L _{n}[{L}_{n}])$.
@@
Recall that $\mathsf{MV}_{n}=\mathbb{Q}(\L _{n})$. Furthermore, $c_{n}\in\mathsf{imp}_{eq}(\L _{n})$ and $\L _{n}[{L}_{n}]$ is defined. Therefore, in view of 13.25, it suffices to show that $\L _{n}[{L}_{n}]$ is a primal algebra. To this end, we will employ 13.23.
@@
(i): Let $\bm{A}\leqslant\bm{B}\in\mathsf{K}[{L}_{\mathcal{F}}]$. To establish the inclusion from left to right, consider $b\in B-\mathsf{d}_{\mathsf{M}}(\bm{A},\bm{B})$. Then there exists $\bm{C}\in\mathsf{M}$ and a pair of homomorphisms $g,h\colon\bm{B}\to\bm{C}$ such that $g{\upharpoonright}_{A}=h{\upharpoonright}_{A}$ and $g(b)\neq h(b)$. From 10.2 it follows that $\bm{C}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{K}$. Together with the fact that $g,h\colon\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}\to\bm{C}{\upharpoonright}_{{L}_{\mathsf{K}}}$ are homomorphisms such that $g{\upharpoonright}_{A}=h{\upharpoonright}_{A}$ and $g(b)\neq h(b)$, this yields $b\in B-\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}})$. Hence, $\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}})\subseteq\mathsf{d}_{\mathsf{M}}(\bm{A},\bm{B})$. To prove the reverse inclusion, consider $b\in B-\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}})$. Then there exist $\bm{C}\in\mathsf{K}$ and a pair of homomorphisms $g,h\colon\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}\to\bm{C}$ such that $g{\upharpoonright}_{A}=h{\upharpoonright}_{A}$ and $g(b)\neq h(b)$. Since $\mathsf{M}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$ is a pp expansion of the universal class $\mathsf{K}$ by assumption, 9.6 guarantees that $\mathsf{K}$ is the class of ${L}_{\mathsf{K}}$-subreducts of $\mathsf{K}[{L}_{\mathcal{F}}]$. So, there exists $\bm{D}\in\mathsf{K}[{L}_{\mathcal{F}}]$ such that $\bm{C}\leqslant\bm{D}{\upharpoonright}_{{L}_{\mathsf{K}}}$. As $\bm{B},\bm{D}\in\mathsf{K}[{L}_{\mathcal{F}}]$, the homomorphisms $g,h\colon\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}\to\bm{D}{\upharpoonright}_{{L}_{\mathsf{K}}}$ can by viewed as homomorphisms $g,h\colon\bm{B}\to\bm{D}$ by 9.5. Therefore, from $\bm{D}\in\mathsf{M}$, $g{\upharpoonright}_{A}=h{\upharpoonright}_{A}$, and $g(b)\neq h(b)$ it follows that $b\in B-\mathsf{d}_{\mathsf{M}}(\bm{A},\bm{B})$. Hence, conclude that $\mathsf{d}_{\mathsf{M}}(\bm{A},\bm{B})\subseteq\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}})$.
@@
(ii): Assume that $\mathsf{M}$ is a Beth companion of $\mathsf{K}$ and consider $\bm{A}\leqslant\bm{B}\in\mathsf{M}$. We will show that $\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}})=A$. Since $\mathsf{M}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$, there exists $\bm{C}\in\mathsf{K}[{L}_{\mathcal{F}}]$ such that $\bm{B}\leqslant\bm{C}$. As $\bm{C}\in\mathsf{K}[{L}_{\mathcal{F}}]$, from (i) it follows that $\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{C}{\upharpoonright}_{{L}_{\mathsf{K}}})=A$. Moreover, 4.6(i) yields $\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}})\subseteq\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{C}{\upharpoonright}_{{L}_{\mathsf{K}}})$. Therefore, $\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}})=A$.
@@
To prove the converse implication assume that $\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}})=A$ for all $\bm{A}\leqslant\bm{B}\in\mathsf{M}$. By 6.6 and 11.6, to show that $\mathsf{M}$ is a Beth companion of $\mathsf{K}$, it suffices to establish that $\mathsf{d}_{\mathsf{M}}(\bm{C},\bm{D})=C$ for all $\bm{C}\leqslant\bm{D}\in\mathsf{M}$. Consider $\bm{C}\leqslant\bm{D}\in\mathsf{M}$. As $\mathsf{M}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$, there exists $\bm{E}\in\mathsf{K}[{L}_{\mathcal{F}}]$ such that $\bm{D}\leqslant\bm{E}$. Then
@@
where the first equality holds by 4.6(i) and the second follows from (i) because $\bm{E}\in\mathsf{K}[{L}_{\mathcal{F}}]$. Our assumption implies that $\mathsf{d}_{\mathsf{K}}(\bm{C}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{E}{\upharpoonright}_{{L}_{\mathsf{K}}})=C$. Thus, $\mathsf{d}_{\mathsf{M}}(\bm{C},\bm{D})=C$. ∎
@@
Since $\mathsf{M}_{1}$ and $\mathsf{M}_{2}$ are Beth companions of $\mathsf{K}$, by 11.7 there exists a pair of maps $\tau\colon{L}_{\mathsf{M}_{1}}\to T_{2}$ and $\rho\colon{L}_{\mathsf{M}_{2}}\to T_{1}$ witnessing that $\mathsf{M}_{1}$ and $\mathsf{M}_{2}$ are faithfully term equivalent relative to $\mathsf{K}$, where $T_{i}$ is the set of terms of $\mathsf{M}_{i}$ in a countably infinite set of variables for $i=1,2$. By symmetry it suffices to show $\mathsf{M}_{1}{\upharpoonright}_{{L}_{\mathsf{K}}}\subseteq\mathsf{M}_{2}{\upharpoonright}_{{L}_{\mathsf{K}}}$. To this end, consider $\bm{A}\in\mathsf{M}_{1}$. The definition of a faithful term equivalence yields $\rho(\bm{A})\in\mathsf{M}_{2}$ and $\rho(\bm{A}){\upharpoonright}_{{L}_{\mathsf{K}}}=\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}$ (see 11.12(iii)). Therefore, $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}=\rho(\bm{A}){\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{M}_{2}{\upharpoonright}_{{L}_{\mathsf{K}}}$, as desired. ∎
We are now ready to prove 13.2.
@@
We first prove the inclusion $\mathsf{K}[{L}_{\mathcal{F}}]{\upharpoonright}_{{L}_{\mathsf{K}}}\subseteq\mathsf{K}_{\textsc{ac}}$. Consider $\bm{A}\in\mathsf{K}[{L}_{\mathcal{F}}]$. To show that $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}$ is absolutely closed in $\mathsf{K}$, let $\bm{B}\in\mathsf{K}$ with $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\leqslant\bm{B}$. We need to prove that $\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{B})=A$. Since $\mathsf{M}$ is a pp expansion of $\mathsf{K}$, 10.2 guarantees the existence of $\bm{C}\in\mathsf{M}$ such that $\bm{B}\leqslant\bm{C}{\upharpoonright}_{{L}_{\mathsf{K}}}$. Then 4.6(i) yields $\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{B})\subseteq\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{C}{\upharpoonright}_{{L}_{\mathsf{K}}})$. As $\mathsf{M}$ is a Beth companion of $\mathsf{K}$, from 13.28(ii) it follows that $\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{C}{\upharpoonright}_{{L}_{\mathsf{K}}})=A$. Therefore, $\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{B})=A$, as desired

### v2

@@ 13. Absolutely closed and primal algebras
Let $\bm{A}$ be an L{L}-algebra, $\mathcal{F}\subseteq\mathsf{imp}_{\textsc{pp}}(\bm{A})$, and ${L}_{\mathcal{F}}$ an $\mathcal{F}$-expansion of L{L} such that $\bm{A}[{L}_{\mathcal{F}}]$ is defined. If $\bm{A}[{L}_{\mathcal{F}}]$ is primal, then $\mathbb{V}(\bm{A}[{L}_{\mathcal{F}}])$ is a Beth companion of $\mathbb{Q}(\bm{A})$. If, moreover, $\mathcal{F}\subseteq\mathsf{imp}_{\textsc{eq}}(\bm{A})$, then the Beth companion $\mathbb{V}(\bm{A}[{L}_{\mathcal{F}}])$ is equational.
@@
Suppose that $\bm{A}[{L}_{\mathcal{F}}]$ is primal. Then $\mathbb{V}(\bm{A}[{L}_{\mathcal{F}}])$ has the strong epimorphism surjectivity property by Proposition 13.24. Furthermore, recall from [21, Thm. IV.9.4] that $\mathbb{Q}(\bm{A}[{L}_{\mathcal{F}}])=\mathbb{V}(\bm{A}[{L}_{\mathcal{F}}])$ because $\bm{A}[{L}_{\mathcal{F}}]$ is primal. Together with Theorem 11.6, this yields that, to conclude that $\mathbb{V}(\bm{A}[{L}_{\mathcal{F}}])$ is a Beth companion of $\mathbb{Q}(\bm{A})$, it only remains to show that $\mathbb{Q}(\bm{A}[{L}_{\mathcal{F}}])$ is a pp expansion of $\mathbb{Q}(\bm{A})$.
Since $\mathcal{F}\subseteq\mathsf{imp}_{\textsc{pp}}(\bm{A})$, for every $f\in\mathcal{F}$ there exists a pp formula $\varphi_{f}$ functional in $\bm{A}$ that defines $f$. By Corollary 3.11 each $\varphi_{f}$ defines some $f^{*}\in\mathsf{imp}_{\textsc{pp}}(\mathbb{Q}(\bm{A}))$. Let $\mathcal{F}^{*}=\{f^{*}:f\in\mathcal{F}\}$. As $\bm{A}[{L}_{\mathcal{F}}]$ is defined, $f^{\bm{A}}$ is total for every $f\in\mathcal{F}$. Consequently, $f^{\bm{A}}=(f^{*})^{\bm{A}}$ because $f$ and $f^{*}$ are both defined by $\varphi_{f}$. Therefore, $(f^{*})^{\bm{A}}$ is total for every $f^{*}\in\mathcal{F}^{*}$. Then Proposition 8.11(ii) yields $\mathcal{F}^{*}\subseteq\mathsf{ext}_{\textsc{pp}}(\mathbb{Q}(\bm{A}))$. We can regard ${L}_{\mathcal{F}}$ as an $\mathcal{F}^{*}$-expansion of L{L} by setting $g_{f^{*}}=g_{f}$ for each $f\in\mathcal{F}$. Since $f^{\bm{A}}=(f^{*})^{\bm{A}}$ for every $f\in\mathcal{F}$, the definition of $\bm{A}[{L}_{\mathcal{F}}]$ is independent on whether ${L}_{\mathcal{F}}$ is thought of as an $\mathcal{F}$-expansion or as an $\mathcal{F}^{*}$-expansion. Thus, Theorem 10.5 implies that $\mathbb{Q}(\bm{A}[{L}_{\mathcal{F}}])$ is a pp expansion of $\mathbb{Q}(\bm{A})$ induced by $\mathcal{F}^{*}$ and ${L}_{\mathcal{F}}$.
@@
For $n\in\mathbb{Z}^{+}$, we denote by $\L _{n}$ the subalgebra of the real unit interval [0,1]{[0,1]} (cf. Example 13.17) with universe $\{\frac{m}{n}:m\in\mathbb{N},\ m\leqslant n\}$. Notice that $\L _{n}$ is a finite $\mathsf{MV}$-algebra of $n+1$ elements. We consider the variety
@@
As $\L _{n}\leqslant{[0,1]}$, Claim 13.19(i, ii) implies that for all $a,b\in\L _{n}$,
@@
Therefore, $\psi_{n}$ defines a total unary $c_{n}\in\mathsf{imp}_{\textsc{eq}}(\L _{n})$ such that $c_{n}^{\scriptsize\L _{n}}$ is the constant map with value $\frac{1}{n}$. Let ${L}_{n}$ be a $c_{n}$-expansion of ${L}_{\mathsf{MV}_{n}}$. Since $c_{n}^{\scriptsize\L _{n}}$ is total, the algebra $\L _{n}[{L}_{n}]$ is defined and can thought of as the result of adding a constant for the element $\frac{1}{n}$ to $\L _{n}$. Then let $\mathsf{DMV}_{n}=\mathbb{V}(\L _{n}[{L}_{n}])$.
@@
Recall that $\mathsf{MV}_{n}=\mathbb{Q}(\L _{n})$. Furthermore, $c_{n}\in\mathsf{imp}_{\textsc{eq}}(\L _{n})$ and $\L _{n}[{L}_{n}]$ is defined. Therefore, in view of Theorem 13.25, it suffices to show that $\L _{n}[{L}_{n}]$ is a primal algebra. To this end, we will employ Theorem 13.23.
@@
(i): Let $\bm{A}\leqslant\bm{B}\in\mathsf{K}[{L}_{\mathcal{F}}]$. To establish the inclusion from left to right, consider $b\in B-\mathsf{d}_{\mathsf{M}}(\bm{A},\bm{B})$. Then there exists $\bm{C}\in\mathsf{M}$ and a pair of homomorphisms $g,h\colon\bm{B}\to\bm{C}$ such that $g{\upharpoonright}_{A}=h{\upharpoonright}_{A}$ and $g(b)\neq h(b)$. From Proposition 10.2 it follows that $\bm{C}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{K}$. Together with the fact that $g,h\colon\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}\to\bm{C}{\upharpoonright}_{{L}_{\mathsf{K}}}$ are homomorphisms such that $g{\upharpoonright}_{A}=h{\upharpoonright}_{A}$ and $g(b)\neq h(b)$, this yields $b\in B-\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}})$. Hence, $\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}})\subseteq\mathsf{d}_{\mathsf{M}}(\bm{A},\bm{B})$. To prove the reverse inclusion, consider $b\in B-\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}})$. Then there exist $\bm{C}\in\mathsf{K}$ and a pair of homomorphisms $g,h\colon\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}\to\bm{C}$ such that $g{\upharpoonright}_{A}=h{\upharpoonright}_{A}$ and $g(b)\neq h(b)$. Since $\mathsf{M}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$ is a pp expansion of the universal class $\mathsf{K}$ by assumption, Proposition 9.6 guarantees that $\mathsf{K}$ is the class of ${L}_{\mathsf{K}}$-subreducts of $\mathsf{K}[{L}_{\mathcal{F}}]$. So, there exists $\bm{D}\in\mathsf{K}[{L}_{\mathcal{F}}]$ such that $\bm{C}\leqslant\bm{D}{\upharpoonright}_{{L}_{\mathsf{K}}}$. As $\bm{B},\bm{D}\in\mathsf{K}[{L}_{\mathcal{F}}]$, the homomorphisms $g,h\colon\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}\to\bm{D}{\upharpoonright}_{{L}_{\mathsf{K}}}$ can by viewed as homomorphisms $g,h\colon\bm{B}\to\bm{D}$ by Proposition 9.5. Therefore, from $\bm{D}\in\mathsf{M}$, $g{\upharpoonright}_{A}=h{\upharpoonright}_{A}$, and $g(b)\neq h(b)$ it follows that $b\in B-\mathsf{d}_{\mathsf{M}}(\bm{A},\bm{B})$. Hence, conclude that $\mathsf{d}_{\mathsf{M}}(\bm{A},\bm{B})\subseteq\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}})$.
@@
(ii): Assume that $\mathsf{M}$ is a Beth companion of $\mathsf{K}$ and consider $\bm{A}\leqslant\bm{B}\in\mathsf{M}$. We will show that $\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}})=A$. Since $\mathsf{M}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$, there exists $\bm{C}\in\mathsf{K}[{L}_{\mathcal{F}}]$ such that $\bm{B}\leqslant\bm{C}$. As $\bm{C}\in\mathsf{K}[{L}_{\mathcal{F}}]$, from (i) it follows that $\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{C}{\upharpoonright}_{{L}_{\mathsf{K}}})=A$. Moreover, Corollary 4.6(i) yields $\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}})\subseteq\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{C}{\upharpoonright}_{{L}_{\mathsf{K}}})$. Therefore, $\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}})=A$.
@@
To prove the converse implication assume that $\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}})=A$ for all $\bm{A}\leqslant\bm{B}\in\mathsf{M}$. By Propositions 6.6 and 11.6, to show that $\mathsf{M}$ is a Beth companion of $\mathsf{K}$, it suffices to establish that $\mathsf{d}_{\mathsf{M}}(\bm{C},\bm{D})=C$ for all $\bm{C}\leqslant\bm{D}\in\mathsf{M}$. Consider $\bm{C}\leqslant\bm{D}\in\mathsf{M}$. As $\mathsf{M}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$, there exists $\bm{E}\in\mathsf{K}[{L}_{\mathcal{F}}]$ such that $\bm{D}\leqslant\bm{E}$. Then
@@
where the first equality holds by Corollary 4.6(i) and the second follows from (i) because $\bm{E}\in\mathsf{K}[{L}_{\mathcal{F}}]$. Our assumption implies that $\mathsf{d}_{\mathsf{K}}(\bm{C}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{E}{\upharpoonright}_{{L}_{\mathsf{K}}})=C$. Thus, $\mathsf{d}_{\mathsf{M}}(\bm{C},\bm{D})=C$. ∎
@@
Since $\mathsf{M}_{1}$ and $\mathsf{M}_{2}$ are Beth companions of $\mathsf{K}$, by Theorem 11.7 there exists a pair of maps $\tau\colon{L}_{\mathsf{M}_{1}}\to T_{2}$ and $\rho\colon{L}_{\mathsf{M}_{2}}\to T_{1}$ witnessing that $\mathsf{M}_{1}$ and $\mathsf{M}_{2}$ are faithfully term equivalent relative to $\mathsf{K}$, where $T_{i}$ is the set of terms of $\mathsf{M}_{i}$ in a countably infinite set of variables for $i=1,2$. By symmetry it suffices to show $\mathsf{M}_{1}{\upharpoonright}_{{L}_{\mathsf{K}}}\subseteq\mathsf{M}_{2}{\upharpoonright}_{{L}_{\mathsf{K}}}$. To this end, consider $\bm{A}\in\mathsf{M}_{1}$. The definition of a faithful term equivalence yields $\rho(\bm{A})\in\mathsf{M}_{2}$ and $\rho(\bm{A}){\upharpoonright}_{{L}_{\mathsf{K}}}=\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}$ (see Remark 11.12(iii)). Therefore, $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}=\rho(\bm{A}){\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{M}_{2}{\upharpoonright}_{{L}_{\mathsf{K}}}$, as desired. ∎
We are now ready to prove Theorem 13.2.
@@
We first prove the inclusion $\mathsf{K}[{L}_{\mathcal{F}}]{\upharpoonright}_{{L}_{\mathsf{K}}}\subseteq\mathsf{K}_{\textsc{ac}}$. Consider $\bm{A}\in\mathsf{K}[{L}_{\mathcal{F}}]$. To show that $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}$ is absolutely closed in $\mathsf{K}$, let $\bm{B}\in\mathsf{K}$ with $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\leqslant\bm{B}$. We need to prove that $\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{B})=A$. Since $\mathsf{M}$ is a pp expansion of $\mathsf{K}$, Proposition 10.2 guarantees the existence of $\bm{C}\in\mathsf{M}$ such that $\bm{B}\leqslant\bm{C}{\upharpoonright}_{{L}_{\mathsf{K}}}$. Then Corollary 4.6(i) yields $\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{B})\subseteq\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{C}{\upharpoonright}_{{L}_{\mathsf{K}}})$. As $\mathsf{M}$ is a Beth companion of $\mathsf{K}$, from Proposition 13.28(ii) it follows that $\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{C}{\upharpoonright}_{{L}_{\mathsf{K}}})=A$. Therefore, $\mathsf{d}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{B})=A$, as desired